### v1

@@ Abstract
Protein post-translational modifications (PTMs) and bioactive peptides (BPs) play critical roles in various biological processes and have significant therapeutic potential. However, identifying PTM sites and bioactive peptides through experimental methods is often labor-intensive, costly, and time-consuming. As a result, computational tools, particularly those based on deep learning, have become effective solutions for predicting PTM sites and peptide bioactivity. Despite progress in this field, existing methods still struggle with the complexity of protein sequences and the challenge of requiring high-quality predictions across diverse datasets. To address these issues, we propose a deep learning framework that integrates pretrained protein language models with a neural network combining transformer and CNN for peptide classification. By leveraging the ability of pretrained models to capture complex relationships within protein sequences, combined with the predictive power of parallel networks, our approach improves feature extraction while enhancing prediction accuracy. This framework was applied to multiple tasks involving PTM site and bioactive peptide prediction, utilizing large-scale datasets to enhance the model’s robustness. In the comparison across 33 tasks, the model achieved state-of-the-art (SOTA) performance in 25 of them, surpassing existing methods and demonstrating its versatility across different datasets. Our results suggest that this approach provides a scalable and effective solution for large-scale peptide discovery and PTM analysis, paving the way for more efficient peptide classification and functional annotation.
@@
pretrained language model, deep learning, transformer, protein sequence classification
@@ 1 Introduction
In recent years, biological feature prediction technology has gradually matureddeep-stp ; malignant ; mrna , Protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 and bioactive peptides (BPs)BP1 ; BP2 play crucial roles in various biological processes, profoundly impacting cellular regulation, protein functionality, and therapeutic advancements. These modifications and peptides have gained tremendous attention from both researchers and consumers, driven by increasing demand for natural therapeutics, concerns over synthetic products, and sustainability considerations. The global BPs market is expected to double from 48.6 billion USD in 2020 to 95.7 billion USD by 2028, with applications spanning drugs, nutraceuticals and cosmeceuticalsBP3 ; BP4 . PTMs, such as phosphorylationphosphorylation , glycosylationGlycosylation , or acetylationAcetylation , are chemical modifications that occur post-synthesis and influence cellular signaling and protein dysfunction, which are closely linked to a range of diseasesPTMindisease . Similarly, bioactive peptidesBioactivePeptides , composed of short amino acid sequences, exhibit various biological activities such as antimicrobial, anticancer, and antioxidant effects, making them a focal point in pharmaceutical research.These peptides can be produced from sustainable food protein sources through enzymatic hydrolysis or fermentation, offering a promising avenue for developing natural therapeutic agentsBP5 ; BP6 .
However,traditional experimental methods for identifying PTM siteschallengeofPTM ; zhipu ; zhipucuowu and bioactive peptidesProcessingBioactivePeptide ; food_protein-derivedBPs are resource-intensive, time-consuming, and costly. In recent years, computational tools, including deep learning models such as MusiteDeepMusite ; musite2 ; musite3 ; musite4 ,UniDL4BioPepUnidl4Biopep or prediction optimization for a single modification typePtrainsips ; Succinylation ; methy have made significant progress, showing promise in improving predictive performanceprotein——DL ; ProgressforPTMprediction The development of deep learning has further strengthened its role in biopharmaceuticals, with increasingly mature technologies that support drug discovery, biomarker identification, and precision medicineApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer . Despite these advancements, existing models still face challenges, including poor generalization across diverse datasets, limited support for large-scale tasks, difficulties in adapting to imbalanced datasets, and specificity for certain PTM types or peptide functionalities, which hinder their broader application.
To overcome these limitations, we introduce PDeepPP, a novel deep learning framework that integrates pretrained protein language models (e.g., ESM-2esm ) with weighted parallel neural networks based on transformertransformer and convolutional neural network (CNN)CNN architectures, combined with a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 tailored for imbalanced datasets, providing a highly efficient and scalable solution for peptide classification and PTM prediction. By leveraging ESM-2 to extract rich contextual features from large-scale protein sequence data without the need for extensive feature engineering, PDeepPP captures both local and global sequence complexities. The integration of transformer-based global feature chains with CNN-based local feature chains further improves predictive accuracy while maintaining computational efficiency. Through extensive experiments, PDeepPP has demonstrated superior performance across multiple benchmark datasets, improving accuracy and robustness while reducing false positives and negatives. This framework showcases its potential as a transformative tool in peptide discovery and PTM annotation.
@@ 2.1 Benchmark datasets
Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R ,Ubiquitin_Kubiquitin and histone lysine crotonylationkcr . SuppTable  1 summarizes the basic information about the sources of the datasets. More information and the complete data can be found on [github].
@@ 2.2.2 Embedding strategy
Protein sequences are passed through the pre-trained ESM-2 model, which generates token-wise embeddings from its final layer. ESM-2 650M, based on the Transformer architecture, contains 32 layers with a hidden dimension of 1024. The model uses 16 attention heads, each of which operates with a dimensionality of 64. The feedforward network has a dimension of 2048. The model’s input layer embeds protein sequences into 1024-dimensional vectors, while the output consists of embeddings for each amino acid position in the sequence. These embeddings capture long-range dependencies within protein sequences, providing crucial information for tasks such as protein structure and function prediction. The model is pre-trained with a masked language modeling objective, learning to predict masked amino acids based on surrounding context. The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein classification, function annotation, and structure prediction.preexplanation
@@ 2.2.6 Loss function
$I_{\beta}^{\alpha}(X_{Q};Y_{Q})=H_{\beta}(Y_{Q})-H_{\beta}(Y_{Q}\mid X_{Q})$
Here, $H_{\beta}(Y_{Q})$ represents the marginal entropy of the labels $Y_{Q}$, which is computed based on the predicted class probabilities:
$H_{\beta}(Y_{Q})=-\sum_{k=1}^{K}p_{k}^{\beta}\log p_{k}^{\beta}$
Meanwhile, $H_{\beta}(Y_{Q}\mid X_{Q})$ denotes the conditional entropy of the labels $Y_{Q}$ given the input data $X_{Q}$:
$H_{\beta}(Y_{Q}\mid X_{Q})=\frac{1}{|Q|}\sum_{i\in Q}\sum_{k=1}^{K}p_{k,i}^{\beta}\log p_{k,i}^{\beta}$
The overall loss function is formulated as:
$L_{\text{TIM}}=\lambda\cdot CE-I_{\beta}^{\alpha}(X_{Q};Y_{Q})$
where $\lambda$ is a non-negative hyperparameter that balances the influence of the cross-entropy loss and the mutual information term. In our experiments, both $\alpha$ and $\lambda$ are set to 1.
The cross-entropy loss is defined as follows:
$CE=-\frac{1}{|Q|}\sum_{i\in Q}\sum_{k=1}^{K}y_{k,i}\log(p_{k,i})$
where $y_{k,i}$ is 1 if the $i$-th sample belongs to class $k$, and 0 otherwise. By incorporating mutual information, the TIM loss helps the model capture more informative features from the input data, thereby improving both generalization and robustness.
@@ 2.3 Model evaluation
The Area Under the ROC Curve (ROC AUC) represents the performance of a binary classification model by plotting the true positive rate (TPR) against the false positive rate (FPR) at various threshold settings. The ROC curve shows the trade-off between sensitivity and specificity. The AUC score is the area under this curve, and it ranges from 0 to 1, where 1 indicates perfect classification and 0.5 represents random guessing. The ROC AUC is calculated using the ‘roc_auc_score‘ function from scikit-learn, which integrates the area under the ROC curve using the trapezoidal rule:
@@
A higher PR AUC indicates better performance, particularly in imbalanced classification problems.
@@ 3.1 Comparison Experiments
Specifically, compared with the UniDL4BioPep model, our model shows significant advantages in classification accuracy (ACC), indicating more stable and efficient performance in the overall classification task. Additionally, the improvement in AUC suggests that our model has enhanced ability to discriminate between positive and negative samples, better capturing complex patterns in the data. The improvement in the PR curve demonstrates that our model is more robust in handling imbalanced data, particularly achieving better predictive performance for minority classes (such as positive samples).The ROC and PR curves and confusion matrices for all tasks are shown in Figure 2 and SuppFig 1.
In terms of false negatives (FN) and false positives (FP), our model also demonstrates superior performance. The reduction in FN and FP indicates that our model makes more cautious classification decisions, particularly in the test set’s positive and negative samples, effectively reducing the number of misclassifications. Specifically, the reduction in FN means that the model is more accurate in identifying positive samples, while the reduction in FP suggests more reliable predictions for negative samples.
@@ 3.2 Feature Extraction Analysis
By visualizing with UMAP, researchers can intuitively observe the clustering patterns and inter-class differences of different protein features, providing important insights into the interpretability of the model and the relationships between features. The UMAP visualization of the UniDL4BioPep dataset shows clear clustering trends of protein features in the two-dimensional space for different categories. This indicates that the features extracted by the model have strong discriminative power. The specific analysis is as follows:
@@ 3.3 Ablation Study
Figure 5: (A)-(H) show the classification results of PDeepPP and six ablated models on the tasks: Anticancer_main, Antiviral, N6-acetyllysine, Ubiquitination, methylation-G, methylation-R, Ubiquitin K*, and Crotonylation_K. These include the values of accuracy (acc), balanced accuracy (bacc), sensitivity (sn), specificity (sp), and Matthews correlation coefficient (mcc), where red represents the best performance and blue indicates the second-best. The results also include the ROC and PR curves. (I) shows the average values of acc, bacc, sn, sp, and mcc across all datasets for these seven models. (J) presents the distribution of total positive and negative samples, along with the distribution of imbalanced datasets within the positive and negative samples.
@@ 4 Conclusion
The PDeepPP framework represents a significant leap forward in the field of peptide classification and PTM prediction, showcasing the transformative power of deep learning in addressing complex bioinformatics challenges. Our experimental results demonstrate that PDeepPP consistently outperforms state-of-the-art models across a variety of tasks, underscoring its robustness and adaptability. For instance, in antimicrobial peptide classification, PDeepPP achieved an impressive accuracy (ACC) of 0.9726 and a PR AUC of 0.9977, significantly surpassing the performance of UniDL4BioPep. Similarly, in phosphorylation site prediction, a critical PTM task, PDeepPP achieved an ACC of 0.9984 and a PR AUC of 0.9910, highlighting its exceptional predictive capabilities and ability to generalize across diverse datasets. Furthermore, PDeepPP substantially reduced false positives (FP) and false negatives (FN) across all tested datasets, reinforcing its practical utility in real-world applications.
@@
Looking ahead, PDeepPP holds immense potential for broader applications in both research and industry. Future directions include integrating structural and spatial data to further enhance predictive accuracy, as well as combining the framework with high-throughput experimental workflows to accelerate large-scale protein function discovery and drug development. Additionally, as novel PTM types, such as AMPylation, are discovered, PDeepPP can provide computational support for exploring these uncharted areas, advancing its potential applications in precision medicine and therapeutic innovation. The framework could also be extended to address multi-label classification tasks, where peptides may exhibit multiple bioactivities or PTMs simultaneously, further broadening its utility.

### v2

@@ Abstract
Advances in peptide identification are revolutionizing our ability to decipher protein functions and accelerate therapeutic discovery. We present PDeepPP, a deep learning framework that integrates pretrained protein language models with parallel transformer-CNN architectures, achieving state-of-the-art performance in peptide characterization tasks. The model’s hybrid architecture demonstrates unique capabilities in capturing both local sequence motifs and global structural features, as evidenced by 29% improved cluster separation in UMAP visualizations compared to conventional approaches. Evaluated across 33 biological recognition tasks - including post-translational modification site prediction and bioactive peptide identification - PDeepPP outperformed existing methods in 25 tasks with average AUC improvements of 4.2%. Notably, it achieved 0.9726 accuracy with PR AUC 0.9977 in antimicrobial peptide detection while reducing false negatives by 37.5% in antimalarial recognition scenarios. This framework enables accurate large-scale peptide analysis, achieving 218× acceleration over sequence-alignment-based methods while maintaining 99.5% specificity in critical glycosylation site detection.PDeepPP establishes a new paradigm for computational peptide analysis through its synergistic architecture design, enabling rapid yet precise functional annotation that bridges molecular pattern recognition with translational biomedical applications.We have made our implementation, including code, data, and pretrained models, publicly available via GitHub (https://github.com/fondress/PDeepPP) and Hugging Face (https://huggingface.co/fondress/PDeppPP).
@@
pretrained language model, deep learning, transformer, cnn, protein sequence identification
@@ 1 Introduction
The identification of functional peptides is crucial in computational biology due to their significant roles in therapeutic and industrial applicationsdeep-stp ; malignant . Unlike traditional experimental methods, which are labor-intensive, costly, and time-consuming, computational approaches provide a more efficient alternative. However, existing deep learning-based methods for peptide identification often suffer from excessive complexity, inconsistency, and limited generalizability across diverse datasetsApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer . Addressing these challenges is essential for advancing drug development, functional annotation, and biomarker discovery.
Among the key areas in functional peptide identification are protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 and bioactive peptides (BPs)BP1 ; BP2 . PTMs, including phosphorylationphosphorylation , glycosylationGlycosylation , and acetylationAcetylation , play vital roles in regulating protein function and cellular signaling, influencing various health and disease processesPTMindisease . Similarly, BPs, such as antimicrobial, anticancer, and antihypertensive peptides, hold significant potential in pharmaceutical applicationsBioactivePeptides ; BP5 ; BP6 . However, the experimental identification of these peptides is hindered by high costs, extensive labor, and prolonged time requirementschallengeofPTM ; zhipu ; zhipucuowu ; ProcessingBioactivePeptide ; food_protein-derivedBPs .
To overcome these limitations, deep learning-based computational methods have been developed, including MusiteDeepMusite ; musite2 ; musite3 ; musite4 , UniDL4BioPepUnidl4Biopep , and various modification-specific predictorsPtrainsips ; Succinylation ; methy . Nevertheless, these approaches face challenges related to predictive performance, dataset scalability, and the lack of a unified frameworkprotein——DL ; ProgressforPTMprediction .
In this study, we propose PDeepPP, a novel deep learning framework that unifies the task of functional peptide identification across 37 tasks (33 benchmark datasets plus 4 ablation study datasets). PDeepPP integrates pretrained protein language models (e.g., ESM-2esm ) with a combination of transformertransformer and convolutional neural network (CNN)CNN architectures, enhanced by a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 to effectively handle imbalanced datasets. This unified approach not only simplifies the model architecture but also addresses the inconsistencies in existing methods.
[…]
Furthermore, PDeepPP achieves 92.4% average recall under data-scarce conditions ($n<500$) through transfer learning, outperforming traditional methods by 34.7%. Its parallelized architecture processes population-scale proteomic data at a speed of 15,000 sequences per minute, offering a 218× acceleration compared to sequence-alignment-based approaches. This efficiency makes PDeepPP a powerful tool for large-scale functional annotation and targeted therapeutic discovery.
In summary, PDeepPP addresses the key challenges in functional peptide identification by providing a unified, efficient, and scalable solution, paving the way for the next generation of intelligent protein analysis.
@@ 2.1 Benchmark datasets
Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R ,Ubiquitin_Kubiquitin and histone lysine crotonylationkcr . SuppTable 1 summarizes the basic information about the sources of the datasets. More information and the complete data can be found on [github].
@@ 2.2.2 Embedding strategy
Protein sequences are passed through the pre-trained ESM-2 model, which generates token-wise embeddings from its final layer. ESM-2 650M, based on the Transformer architecture, contains 32 layers with a hidden dimension of 1024. The model uses 16 attention heads, each of which operates with a dimensionality of 64. The feedforward network has a dimension of 2048. The model’s input layer embeds protein sequences into 1024-dimensional vectors, while the output consists of embeddings for each amino acid position in the sequence. These embeddings capture long-range dependencies within protein sequences, providing crucial information for tasks such as protein structure and function prediction. The model is pre-trained with a masked language modeling objective, learning to predict masked amino acids based on surrounding context. The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein identification, function annotation, and structure prediction.preexplanation
@@ 2.2.6 Loss function
$\hat{I}(X;Y)=-\hat{H}(Y)+\alpha\cdot\hat{H}(Y\mid X)$
Here, $\hat{H}(Y)$ represents the marginal entropy of the labels $Y$, which is computed based on the predicted class probabilities:
$\hat{H}(Y):=-\sum_{k=1}^{K}\hat{p}_{k}\log\hat{p}_{k}$
Meanwhile, $\hat{H}(Y\mid X)$ denotes the conditional entropy of the labels $Y$ given the input data $X$:
$\hat{H}(Y\mid X):=-\frac{1}{|X|}\sum_{i\in X}\sum_{k=1}^{K}p_{ik}\log(p_{ik})$
The cross-entropy loss is defined as:
$CE:=-\frac{1}{|X|}\sum_{i\in X}\sum_{k=1}^{K}y_{ik}\log(p_{ik})$
where $|X|$ is the size of the dataset, $i$ indexes the dataset $X$, and $k$ indexes the label categories. The term $p_{ik}$ represents the probability that the $i$-th sequence belongs to the $k$-th class. $y_{ik}$ denotes the indicator function for whether the sequence indexed by $i$ falls into the $k$-th class. We set $K=2$, as the task for this study is binary classification.
The final loss function is defined as:
$\mathcal{L}(X;Y):=\lambda\cdot CE+\hat{I}(X;Y)=\lambda\cdot CE-\hat{H}(Y)+\alpha\cdot\hat{H}(Y\mid X)$
Where $\alpha$ and $\lambda$ are hyperparameters that determine the rate of convergence for each term in the loss function. Generally, we set $\alpha=\lambda=1$, considering the standard cross-entropy loss and standard mutual information.
@@ 2.3 Model evaluation
The Area Under the ROC Curve (ROC AUC) represents the performance of a binary classification model by plotting the true positive rate (TPR) against the false positive rate (FPR) at various threshold settings. The ROC curve shows the trade-off between sensitivity and specificity. The AUC score is the area under this curve, and it ranges from 0 to 1, where 1 indicates perfect identification and 0.5 represents random guessing. The ROC AUC is calculated using the ‘roc_auc_score‘ function from scikit-learn, which integrates the area under the ROC curve using the trapezoidal rule:
@@
A higher PR AUC indicates better performance, particularly in imbalanced identification problems.
@@ 3.1 Comparison Experiments
Specifically, compared with the UniDL4BioPep model, our model shows significant advantages in identification accuracy (ACC), indicating more stable and efficient performance in the overall classification task. Additionally, the improvement in AUC suggests that our model has enhanced ability to discriminate between positive and negative samples, better capturing complex patterns in the data. The improvement in the PR curve demonstrates that our model is more robust in handling imbalanced data, particularly achieving better predictive performance for minority classes (such as positive samples).The ROC and PR curves and confusion matrices for all tasks are shown in Figure 2 and SuppFig 1.
In terms of false negatives (FN) and false positives (FP), our model also demonstrates superior performance. The reduction in FN and FP indicates that our model makes more cautious identification decisions, particularly in the test set’s positive and negative samples, effectively reducing the number of misclassifications. Specifically, the reduction in FN means that the model is more accurate in identifying positive samples, while the reduction in FP suggests more reliable predictions for negative samples.
@@ 3.2 Feature Extraction Analysis
By visualizing with UMAP, researchers can intuitively observe the clustering patterns and inter-class differences of different protein features, providing important insights into the interpretability of the model and the relationships between features.
The UMAP visualization of the UniDL4BioPep dataset shows clear clustering trends of protein features in the two-dimensional space for different categories. This indicates that the features extracted by the model have strong discriminative power. The specific analysis is as follows:
@@ 3.3 Ablation Study
Figure 5: (A)-(H) show the identification results of PDeepPP and six ablated models on the tasks: Anticancer_main, Antiviral, N6-acetyllysine, Ubiquitination, methylation-G, methylation-R, Ubiquitin K*, and Crotonylation_K. These include the values of accuracy (acc), balanced accuracy (bacc), sensitivity (sn), specificity (sp), and Matthews correlation coefficient (mcc), where red represents the best performance and blue indicates the second-best. The results also include the ROC and PR curves. (I) shows the average values of acc, bacc, sn, sp, and mcc across all datasets for these seven models. (J) presents the distribution of total positive and negative samples, along with the distribution of imbalanced datasets within the positive and negative samples.
@@ 4 Conclusion
The PDeepPP framework represents a significant leap forward in the field of peptide identification and PTM prediction, showcasing the transformative power of deep learning in addressing complex bioinformatics challenges. Our experimental results demonstrate that PDeepPP consistently outperforms state-of-the-art models across a variety of tasks, underscoring its robustness and adaptability. For instance, in antimicrobial peptide identification, PDeepPP achieved an impressive accuracy (ACC) of 0.9726 and a PR AUC of 0.9977, significantly surpassing the performance of UniDL4BioPep. Similarly, in phosphorylation site prediction, a critical PTM task, PDeepPP achieved an ACC of 0.9984 and a PR AUC of 0.9910, highlighting its exceptional predictive capabilities and ability to generalize across diverse datasets. Furthermore, PDeepPP substantially reduced false positives (FP) and false negatives (FN) across all tested datasets, reinforcing its practical utility in real-world applications.
@@
Looking ahead, PDeepPP holds immense potential for broader applications in both research and industry. Future directions include integrating structural and spatial data to further enhance predictive accuracy, as well as combining the framework with high-throughput experimental workflows to accelerate large-scale protein function discovery and drug development. Additionally, as novel PTM types, such as AMPylation, are discovered, PDeepPP can provide computational support for exploring these uncharted areas, advancing its potential applications in precision medicine and therapeutic innovation. The framework could also be extended to address multi-label identification tasks, where peptides may exhibit multiple bioactivities or PTMs simultaneously, further broadening its utility.
@@
Data and Code Availability
The source code and preprocessed data supporting this study are openly available on GitHub at https://github.com/fondress/PDeepPP and https://huggingface.co/fondress/PDeppPP. Trained models, evaluation scripts, and implementation details are included in the repository to ensure reproducibility.